OpenStudy (anonymous):

For which value of x does the graph of y = 3x^2 - 2x - 5 cross the x-axis?

OpenStudy (callisto):

y = 3x^2 - 2x - 5 y = (3x-5)(x+1) When it crosses x-axis, y=0, so put y=0 into the equation 0 = (3x-5)(x+1) (3x-5)=0 or (x+1)=0 Can you solve them?

OpenStudy (anonymous):

Just follow Callisto's steps and solve for x: 3x - 5 = 0 & x + 1 = 0

OpenStudy (anonymous):

-1?

OpenStudy (anonymous):

Yes that's the second one.

OpenStudy (callisto):

-1 is one of the answers. You also need to solve (3x-5)=0

OpenStudy (anonymous):

im confused on that 1 :S

OpenStudy (callisto):

3x-5 =0 3x-5 +5 = 0+5 3x = 5 \[\frac{3x}{3} = \frac{5}{3}\] x=?

OpenStudy (anonymous):

-2?

OpenStudy (callisto):

Nope :| \[x=\frac{5}{3}\] :|